A 2.44 wide channel with a bed slope of 0.0004 has a depth of flow of 0.61 m.
(a) Determine the discharge of the channel. The manning coefficient is 0.015.
(b) Determine the normal depth if the channel has a flow rate of 2.85 m3/s.
(c) Determine the cross section of greatest hydraulic efficiency for a trapezoidal channel if the design discharge is 10 m3/s, the channel slope is 0.0052 and manning’s n = 0.025.
Answer #1

Y = 0.610 m O. 015 Area = Bxy - 2.44 x 0.16 Perim etu P = B+2y Ā ?0.3904m Hy draulie Radtiug (R] > - 2-u4 + 2%0.6) = 3.c0 m 0Slope = 0.052 n: 0.025 d=60° Hydraulic efficient Trapcaoidal channel -b most ef ficient 29155 TYapezoided channel ) half 9/13